OpenStudy (anonymous):

Julie and Brian are financing $245,000 to purchase a house. They obtained a 30/5 balloon mortgage at 5.25%. What will their balloon payment be?

OpenStudy (amistre64):

in effect, you need to determine the monthly payments for a 30 year fixed loan, and balance it out oafter 5 years

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

here are the choices $223,884.91 $226,753.98 $229,679.66 $234,068.73

OpenStudy (amistre64):

then we have something to compare our results with .. all we do now is work the process.

OpenStudy (anonymous):

i think its the 3rd one

OpenStudy (amistre64):

5 years out of 30 aint a lot ... my gut is on the 4th one

OpenStudy (anonymous):

why do you think

OpenStudy (amistre64):

becuase 5 years out of 30 ... i thought i posted that already :/

OpenStudy (anonymous):

oh i see it now

OpenStudy (amistre64):

lets work the process and make sure :)

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

i have my own formula i use to determine these things which isnt a textbook formula ... what formulas do you ahve to determine monthly payments, and the balance in an account after so many years?

OpenStudy (anonymous):

could i try yours

OpenStudy (amistre64):

you could ... i like it better than the texts

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

the balance in an account is the result of taking the original balance compounded with interest and adjusting it for a geometric progression of payments ... \[A=Bk^n-P\frac{1-k^n}{1-k}\] where, k=1+r/12 to determine payments, A=0 when the loan is paid off, so we can solve for P \[P=Bk^n\frac{1-k}{1-k^n}\] since its a 30 year loan ... n=30*12

OpenStudy (anonymous):

ohhhh

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=245000k%5E%2830*12%29+%281-k%29%2F%281-k%5E%2830*12%29%29%2C+k%3D1%2B.0525%2F12 so we get monthly payments of 1352.90 to play with, and we can then determine the value of A, the account balance, after 5 years using that payment value

OpenStudy (amistre64):

\[A=245,000~k^{5*12}-P\frac{1-k^{5*12}}{1-k}\]

OpenStudy (anonymous):

ok i see i think i can continue on my own now! thank you so much!

OpenStudy (amistre64):

yep ... i gotta rechk my inputs to make sure i didnt type in a bad value ... but thats how i would process this question.

OpenStudy (anonymous):

i see thanks!

OpenStudy (amistre64):

yeah, it looks good to me, but i keep getting something of about 225767 which isnt an option. which makes me have to wonder

OpenStudy (amistre64):

at any rate, my guess was off :)